OpenStudy (anonymous):

Find the sum and product of the roots. 2x^2 + 6x - 8 = 0

OpenStudy (asadkarim7):

2x^2 - 2x +8x -8 = 0 2x ( x - 1 ) + 8 ( x - 1) = 0 2x-8 = 0 ; x-1= 0 x= 4 ; or 1

OpenStudy (anonymous):

u need not do the full solution For a quadratic equation, - (Coefficient of x) - (b) - (6) Sum of roots = --------------------- = -------- = ------- = -3 Coefficient of x^2 a 2 Constant Term c -8 Product of roots = ------------------ = ------- = ------- = -4 Coefficient of x^2 a 2

OpenStudy (anonymous):

you do not need to solve for the roots to answer this question, although the above answers are correct. you have \[x^2+3x-4=0\] so produce of two roots is -4 and sum of two roots is -3

OpenStudy (anonymous):

Yes, asadkarim made a small mistake in his calculation, which is why those don't agree with the above.